3. Comparative advantage and gains from trade Consider two neighboring island countries called Arcadia and Felicidad. They each have 4 million labor hours available per week that they can use to produce jeans, corn, or a combination of both. The following table shows the amount of jeans or corn that can be produced using 1 hour of labor Jeans Corn Country Pairs per hour of labor) (Bushels per hour of labor) Arcadia Felicidad 12 24 8 32 Initially, suppose Arcadia uses 1 million hours of labor per week to produce jeans and 3 million hours per week to produce corn, while Felicidad uses 3 million hours of labor per week to produce jeans and 1 million hours per week to produce corn. Consequently, Arcadia produces 12 million pairs of jeans and 72 million bushels of corn, and Felicidad produces 24 million pairs of jeans and 32 millon bushels of corn. Assume there are no other countries willing to trade goods, so in the absence of trade between these two countries, each country consumes the amount of jeans and corn it produces Arcadias opportunity cost of producing 1 pair of jeans is 1/2 bushel ▼ or corn, and Felicidads opportunity cost of producing 1 pair of Jeans is of corn. Therefore has a comparative advantage in the production of jeans, and has a comparative advantage in the production of corn Suppose that each country completely specializes in the production of the good in which it has a comparative advantage, producing only that good. In this case, the country that produces jeans will produce million pairs per week, and the country that produces corn will produce million bushels per weekArcadia Felicidad Jeans Corn Jeans Corn (Millions of pairs) (Millions of bushels) (Millions of pairs) (Millions of bushels) Without Trade Production 12 72 24 32 Consumption 12 72 24 32 With Trade Production Imports/Exports Consumption Gains from Trade Increase in Consumption Suppose the country that produces jeans trades 26 million pairs of jeans to the other country in exchange for 78 million bushels of corn In the previous table, use the dropdown menus across the row labeled Imports/Exports to select the amount of each good that each country imports and exports. Then enter each countrys final consumption of each good on the line labeled Consumption. When the two countries did not specialize, the total production of jeans was 36 million pairs per week, and the total production of corn was 104 million bushels per week. Because of specialization, the total production of jeans has increased by corn has increased by million pairs per week, and the total production of million bushels per week Because the two countries produce more jeans and more corn under specialization, each country is able to gain from trade Calculate the gains from trade-that is, the amount by which each country has increased its consumption of each good relative to the first row of the previous table. Enter this difference in the boxes across the last row (labeled Increase in Consumption)
Answer #1

Arcadias opportunity cost of producing 1 pair of jeans is 24/12 = 2 bushels of corn and Felicidad's opportunity cost of producing 1 pair of jeans is 32/8=4 bushels of corn. Because Arcadia has a lower opportunity cost for the production of jeans. Therefore, Arcadia has a comparative advantage in the production of jeans and Felicidad has a comparative advantage in the production of corn.

Suppose that each country completely specializes in the production of the good in which it has a comparative advantage , producing only that good . In this case,the country that produces jeans i.e Arcadia will produce (4)(12) = 48 million pairs per week And the country that produces corn  i.e Felicidad will produce 4(32)=128 million bushels per week.

Suppose the country that produces jeans  trades 26 million pairs of jeans to the other country in exchange for 78 million bushels of corn.

ARCADIA FELICIDAD
Jeans (milions of pairs) Corn (Millions of bushels) Jeans (Millions of pairs) Corn (Millions of bushels)
Without trade Production 12 72 24 32
Consumption 12 72 24 32
With Trade Production 48 0 0 128
Trade action Export 26 Import 78 Import 26 Export 78
Consumption (48-26)=22 78 26 (128-78)=50
Gains from trade Increase in consumption (22-12)=10 (78-72)=6 (26-24)=2 (50-32)=18

When the two countries did not specialize ,the total production of jeans was 36 million pairs per week and the total production of corn was 104 million bushels per week.. Because of specialization ,the total production of jeans was 48 million pairs per week and total production of corn was 128 million bushels  per week. This implies that because of specialization , total production of jeans has increased by (48-36)=12 million pairs per week and total production of corn  has increased by (128-104)= 24 million bushels per week.